OpenStudy (anonymous):

help pweaseeeeeeeeeeee You put $600 in a savings account with an APR of 4.2% that compounds quarterly. Assume you do not make any other deposits or withdrawals a. How much money is in the account after 3 years? Either copy and paste your table or show your calculation. b. About how long will it take for your money to double?

OpenStudy (anonymous):

@phi @amistre64 @unklerhaukus @cwrw238

OpenStudy (amistre64):

is APR the same as interest? or is that something else?

OpenStudy (anonymous):

yea annual percentage rate

OpenStudy (amistre64):

are you spose to do this recursively? or can you just use the formula for compounding interest?

OpenStudy (anonymous):

i needa know if my answers are right i got 1 625.5997 2 652.2916 3 680.1224

OpenStudy (anonymous):

and for second question i got 16-17 years

OpenStudy (amistre64):

your 123s look fine to me

OpenStudy (amistre64):

might wanna round to them tho

OpenStudy (amistre64):

and yeah, it takes over 16 but less than 17 years

OpenStudy (amistre64):

if you change the time to months, instead of years, you can prolly get a more exact time

OpenStudy (amistre64):

or A=P(1+r/n)^(nt) A/P = (1+r/n)^(nt) log_{1+r/n}(A/P) = nt (log_{1+r/n}(A/P))/n = t (ln(A/P))/(n ln_{1+r/n}) = t (ln(2))/(4 ln(1+.042/4)) = t = 16.59 years